Can someone PLEASE help me with this problem? THANK YOU!
Let z1 = r1(cos 1 + isin 1) and z2 = r2(cos 2 + isin 2) be two complex
numbers. Show that if z2 (doesn't equal to) 0, then (z1/z2)=(r1/r2)[cos(1 2) + isin(1 2)].

Answer by stanbon(75887)

Let z1 = r1(cos 1 + isin 1) and z2 = r2(cos 2 + isin 2) be two complex
numbers. Show that if z2 (doesn't equal to) 0, then (z1/z2)=(r1/r2)[cos(1 2) + isin(1 2)].
---
z1/z2 = [r1cos(1) + r1isin(1)]/[r2cos(2) + r2isin(2)]
------
Multiply numerator and denominator by [r2cos(2)-r2isin(2)] to get:
-------
z1/z2
= [[r1cos(1) + r1isin(1)][r2cos(2)-r2isin(2)]]/[(r2cos(2))^2 + (r2sin(2))^2]]
----
I think I'll let you continue this.
=======================================
Cheers,
Stan H.